OpenStudy (anonymous):

Eliminate the parameter. x = t^2 + 2, y = t^2 - 4

OpenStudy (anonymous):

I think the answer is y = x - 6, x ≥ 1.

OpenStudy (aaronq):

take one of the equations and isolate t, sub the expression you get into the other equation and simplify.

OpenStudy (anonymous):

So I'm at √x-2 = 6?

OpenStudy (aaronq):

hm you took the one on the right? you should have \( t=\sqrt{x -2}\) then \( y = t^2 - 4 \rightarrow y = (\sqrt{x -2})^2 - 4 \)

OpenStudy (anonymous):

Wait the six should be t, lol sorry. And okay then what?